Solve the following differential equations:
i) y sin2x dx = (1+ y²+ cos²x) dy
ii) (xy² - x²)dx + (3x²y² + x²y - 2x³ + y²)dy = 0
1
Expert's answer
2020-03-23T12:25:04-0400

i)

"ysin2xdx-(1+y^2+cos^2x)dy=0\\
P(x,y)=ysin2x, \\
Q(x,y)=-(1+y^2+cos^2x)\\
\frac{\partial P}{\partial y}=sin2x\\
\frac{\partial Q}{\partial x}=-2cosx\cdot (-sinx)=sin2x\\
\frac{\partial P}{\partial y}=\frac{\partial Q}{\partial x}"

equation in total differentials

"\exist u(x,y): \frac{\partial u}{\partial x}=P(x,y)\\
\frac{\partial u}{\partial y}=Q(x,y)\\"

In our case

"\frac{\partial u}{\partial x}=ysin2x\\
\frac{\partial u}{\partial y}=-(1+y^2+cos^2x)\\
u(x,y)=\int ysin2xdx=-\frac{y}{2}cos2x+\phi(y)"

Find

"\frac{\partial u}{\partial y}=-\frac{1}{2}cos2x+\phi'(y)=\\
=-\frac{1}{2}(2cos^2x-1)+\phi'(y)=\frac{1}{2}-cos^2x+\phi'(y)"

Then

"\frac{1}{2}-cos^2x+\phi'(y)=-1-y^2-cos^2x\\
\phi'(y)=-\frac{3}{2}-y^2\\
\phi(y)=-\frac{3}{2}y-\frac{y^3}{3}\\
u(x,y)=-\frac{y}{2}cos2x-\frac{3}{2}y-\frac{y^3}{3}"

Answer: solution of equation is

"-\frac{y}{2}cos2x-\frac{3}{2}y-\frac{y^3}{3}=C"


ii)

"(xy^2-x^2)dx+(3x^2y^2+x^2y-2x^3+y^2)dy=0\\
P(x,y)=xy^2-x^2\\
Q(x,y)=3x^2y^2+x^2y-2x^3+y^2\\
\frac{\partial P}{\partial y}=2xy\\
\frac{\partial Q}{\partial x}=6xy^2+2xy-6x^2\\
\frac{\partial P}{\partial y}\neq\frac{\partial Q}{\partial x}"

Find function "\mu(y)"

"\mu(y)=e^{\int F(y)dy}"

where

"F(y)=\frac{\frac{\partial P}{\partial y}-\frac{\partial Q}{\partial x}}{-P}=\\
=\frac{2xy-6xy^2-2xy+6x^2}{x^2-xy^2}=\frac{6(x^2-xy^2)}{x^2-xy^2}=6\\
\mu(y)=e^{\int6dy}=e^{6y}"

we multiply the equation by "e^{6y}"

"(xy^2-x^2)e^{6y}dx+\\
+(3x^2y^2+x^2y-2x^3+y^2)e^{6y}dy=0\\
P_1(x,y)=(xy^2-x^2)e^{6y}\\
Q_1(x,y)=(3x^2y^2+x^2y-2x^3+y^2)e^{6y}\\
\frac{\partial P_1}{\partial y}=2xye^{6y}+6e^{6y}(xy^2-x^2)=\\
=e^{6y}(2xy+6xy^2-6x^2)\\
\frac{\partial Q_1}{\partial x}=(6xy^2+2xy-6x^2)e^{6y}\\
\frac{\partial P_1}{\partial y}=\frac{\partial Q_1}{\partial x}"

equation in total differentials

"\exist u(x,y): \frac{\partial u}{\partial x}=P_1(x,y)\\
\frac{\partial u}{\partial y}=Q_1(x,y)\\"

in our case

"\frac{\partial u}{\partial x}=(xy^2-x^2)e^{6y}\\
\frac{\partial u}{\partial y}=(3x^2y^2+x^2y-2x^3+y^2)e^{6y}\\
u=\int((xy^2-x^2)e^{6y})dx=(\frac{x^2y^2}{2}-\frac{x^3}{3})e^{6y}+\phi(y)"

Find

"\frac{\partial u}{\partial y}=x^2ye^{6y}+(\frac{x^2y^2}{2}-\frac{x^3}{3})6e^{6y}+\phi'(y)"

then

"x^2ye^{6y}+(\frac{x^2y^2}{2}-\frac{x^3}{3})6e^{6y}+\phi'(y)=\\
=(3x^2y^2+x^2y-2x^3+y^2)e^{6y}\\
\phi'(y)=y^2e^{6y}\\
\phi(y)=\int y^2e^{6y}dy=\frac{1}{6}\int y^2de^{6y}=\\
=\frac{1}{6}(y^2e^{6y}-\int 2ye^{6y}dy)=\frac{1}{6}(y^2e^{6y}-\frac{1}{3}\int yde^{6y})=\\
=\frac{1}{6}(y^2e^{6y}-\frac{1}{3}(ye^{6y}-\int e^{6y}dy))=\\
=\frac{1}{6}y^2e^{6y}-\frac{1}{18}ye^{6y}+\frac{1}{108}e^{6y}"

"u=(\frac{x^2y^2}{2}-\frac{x^3}{3})e^{6y}+\frac{1}{6}y^2e^{6y}-\frac{1}{18}ye^{6y}+\frac{1}{108}e^{6y}"

Answer: solution of equation is

"(\frac{x^2y^2}{2}-\frac{x^3}{3})e^{6y}+\frac{1}{6}y^2e^{6y}-\frac{1}{18}ye^{6y}+\frac{1}{108}e^{6y}=c"